(3,5);8x+y=7 Write an equation. The equation ofn the lineis y=
Answer by mananth(16946)

(3,5);8x+y=7
add -8x to both sides
8x+y-8x=-8x+7
y= -8x+7
this is of the form y= mx +b where m is the slope and b the intercept
m=-8
..
the line passes through (3,5)
5=(-8)*(3)+b
5=-24+b
b=29
5+24=b
so the equation will be
y=-8x+29